### v3

# BSRBF-KAN: A combination of B-splines and Radial Basis Functions in Kolmogorov-Arnold Networks

Abstract
In this paper, we introduce BSRBF-KAN, a Kolmogorov Arnold Network (KAN) that combines B-splines and radial basis functions (RBFs) to fit input vectors during data training. We perform experiments with BSRBF-KAN, multi-layer perception (MLP), and other popular KANs, including EfficientKAN, FastKAN, FasterKAN, and GottliebKAN over the MNIST and Fashion-MNIST datasets. BSRBF-KAN shows stability in 5 training runs with a competitive average accuracy of 97.55% on MNIST and 89.33% on Fashion-MNIST and obtains convergence better than other networks. We expect BSRBF-KAN to open many combinations of mathematical functions to design KANs. Our repo is publicly available at: https://github.com/hoangthangta/BSRBF_KAN.
Keywords Kolmogorov Arnold Networks  $\cdot$ B-splines  $\cdot$ Radial Basis Functions  $\cdot$ Gaussian RBFs  $\cdot$ MNIST
1 Introduction
A recent work of  Liu et al. [1] mentioned applying learnable activation functions as "edges" to fit training data instead of using fixed ones as "nodes" that are usually used in multi-layer perceptions (MLPs). The theory behind KANs relies on the Kolmogorov-Arnold representation theorem (KART), which states that a continuous function of multiple variables can be expressed as a combination of continuous functions of a single variable through additions [2]. KANs are anticipated to bring a fresh perspective to solving issues that have been overshadowed by MLPs.
With that inspiration, many scientists have flocked to develop different types of KANs based on popular polynomials and basis functions. While these works focus on univariate functions to set up KANs, none have explored their combination. Therefore, we aim to combine B-splines [3] and RBFs [4] to build a combined KAN named BSRBF-KAN. The reason why we use them is their successful implementation in two popular KANs, EfficientKAN [^1] https://github.com/Blealtan/efficient-kan and FastKAN [5]. Additionally, they are continuous basis functions used for function approximation, providing features such as smoothness, adaptability, and local control. With the experiments on MNIST [6] and Fashion-MNIST [7], we show the stability and faster convergence of BSRBF-KAN compared to other networks. In this paper, our primary contribution is the development of a combined KAN, BSRBF-KAN, which we compare with other networks on two popular datasets (MNIST and Fashion-MNIST) to understand our KAN behaviors in the image classification problem.
The rest of this document is structured as follows: Section 2 contains related works on KART and KANs. Section 3 outlines our methodology, describing the KART and the design of the KAN architecture, as well as several currently popular KANs. Section 4 presents our experiments, where we compare BSRBF-KAN with MLP and other KANs using results from the MNIST and Fashion-MNIST datasets. This section also includes ablation studies to identify the key components of BSRBF-KAN. Finally, Section 5 provides conclusions, comments, and our future work.
2 Related Works
In 1957, Kolmogorov provided proof to solve Hilbert’s 13th problem, demonstrating that a multivariate continuous function can be represented as a combination of single-variable functions and additions, known as KART [2, 8]. This theorem has been applied in numerous works to build neural networks [9, 10, 11, 12, 13, 14]. However, it has also sparked debates about KART’s ability in designing neural networks. Girosi and Poggio [15] indicated that KART is irrelevant to neural networks because the inner function $\phi_{q,p}$ in Equation 1 can be highly non-smooth [16], preventing $f$ from being smooth, which is essential for generalization and noise stability in neural networks. In contrast, Kůrková [17] argued that KART is relevant to neural networks, demonstrating that all one-variable functions can be accurately approximated by linear combinations of affine functions with some sigmoidal functions. This suggests that while KART may sometimes be unsuitable for neural networks due to its potential non-smooth functions, there are cases where it can be effectively used if appropriate functions are identified.
Despite the long history of application in neural networks, no works on KART had gained significant attention within the research community until the recent work by Liu et al. [1]. They recommended not adhering strictly to the KART but generalizing it to create KANs with more neurons and layers. Our intuition thinks this is an important point to avoid the non-smooth problem when applying KART to neural networks. As a result, KANs can surpass MLPs in both accuracy and interpretability on small-scale AI + Science tasks. However, the revival of KAN has also faced criticism by Dhiman [18], where KANs are described as MLPs with spline-based activation functions, unlike traditional MLPs, which have fixed activation functions. This may offer insights that enhance our understanding of KANs and their comparison with MLPs.
KANs are a fresh breeze that has blown into the scientific community since May 2024. Many KAN design works rely on popular mathematical functions, especially ones that can deal with curves such as B-Splines [3] (Original KAN [1], EfficientKAN [^2] https://github.com/Blealtan/efficient-kan), Gaussian Radial Basis Functions or GRBFs (FastKAN [5]), Reflection SWitch Activation Function or RSWAF (FasterKAN [19]), Chebyshev polynomials (TorchKAN [20], Chebyshev KAN [21]), Legendre polynomials (TorchKAN [20]), Fourier transform (FourierKAN [^3] https://github.com/GistNoesis/FourierKAN/, FourierKAN-GCF [22]) and other polynomial functions [23]. In this paper, we prefer to work on combining basic mathematical functions to design KANs instead of single functions to explore the effects of combined functions on KANs.
3 Methodology
3.1 Kolmogorov-Arnold Representation Theorem
A KAN relies on KART, which states that any multivariate continuous function $f$ defined on a bounded domain can be expressed using a finite number of continuous single-variable functions and additions [24, 25]. Given $\mathbf{x}={x_{1},x_{2},..,x_{n}}$ consisting of $n$ variables, a multivariate continuous function $f(\mathbf{x})$ is represented by:
$f(\mathbf{x})=f(x_{1},\ldots,x_{n})=\sum_{q=1}^{2n+1}\Phi_{q}\left(\sum_{p=1}^{n}\phi_{q,p}(x_{p})\right)$ (1)
which contains two sums: outer sum and inner sum. The outer sum $\sum_{q=1}^{2n+1}$ will sum $12n+$ 1 of $\Phi_{q}$ ($\mathbb{R}\to\mathbb{R}$) terms. The inner one sums up $n$ terms for each $q$, and each term $\phi_{q,p}$ ($\phi_{q,p}\colon[0,1]\to\mathbb{R}$) represents a continuous function of a single variable $x_{p}$.
3.2 The design of KANs
Let’s revisit an MLP characterized by affine transformations and non-linear functions. From an input $\mathbf{x}$, the network performs the composition of weight matrices by layers (from layer 00 to layer $L-1$) and the non-linearity $\sigma$ to generate the final output.
$\text{MLP}(\mathbf{x})$ $=(W_{L-1}\circ\sigma\circ W_{L-2}\circ\sigma\circ\cdots\circ W_{1}\circ\sigma\circ W_{0})\mathbf{x}$ (2)
$=\sigma\left(W_{L-1}\sigma\left(W_{L-2}\sigma\left(\cdots\sigma\left(W_{1}\sigma\left(W_{0}\mathbf{x}\right)\right)\right)\right)\right)$
Liu et al. [1] was inspired by KART to design KANs, but the authors advised generalizing it to more widths and depths. In Equation 1, we must search proper $\Phi_{q}$ and $\phi_{q,p}$ to solve the problem. A general KAN network consisting of $L$ layers takes $\mathbf{x}$ to generate the output as:
$\text{KAN}(\mathbf{x})=(\Phi_{L-1}\circ\Phi_{L-2}\circ\cdots\circ\Phi_{1}\circ\Phi_{0})\mathbf{x}$ (3)
which $\Phi_{l}$ is the function matrix of the $l^{th}$ KAN layer or a set of pre-activations. Let denote the neuron $i^{th}$ of the layer $l^{th}$ and the neuron $j^{th}$ of the layer $l+1^{th}$. The activation function $\phi_{l,i,j}$ connects $(l,i)$ to $(l+1,j)$:
$\phi_{l,j,i},\quad l=0,\cdots,L-1,\quad i=1,\cdots,n_{l},\quad j=1,\cdots,n_{l+1}$ (4)
with $n_{l}$ is the number of nodes of the layer $l^{th}$. So now, the function matrix $\Phi_{l}$ can be represented as a matrix $n_{l+1}\times n_{l}$ of activations:
$\Phi_{l}=\begin{pmatrix}\phi_{l,1,1}(\cdot)&\phi_{l,1,2}(\cdot)&\cdots&\phi_{l,1,n_{l}}(\cdot)\\ \phi_{l,2,1}(\cdot)&\phi_{l,2,2}(\cdot)&\cdots&\phi_{l,2,n_{l}}(\cdot)\\ \vdots&\vdots&\vdots&\vdots\\ \phi_{l,n_{l+1},1}(\cdot)&\phi_{l,n_{l+1},2}(\cdot)&\cdots&\phi_{l,n_{l+1},n_{l}}(\cdot)\end{pmatrix}$ (5)
3.3 Implementation of the current KANs
Original KAN was implemented by Liu et al. [1] by using the residual activation function $\phi(x)$ as the sum of the base function and the spline function with their corresponding weight matrices $w_{b}$ and $w_{s}$:
$\phi(x)=w_{b}b(x)+w_{s}spline(x)$ (6)
where $b(x)$ equals to $silu(x)$ and $s(x)$ is expressed as a linear combination of B-splines. Each activation function is activated with $w_{s}=1$ and $spline(x)\approx 0$, while $w_{b}$ is initialized by using Xavier initialization.
EfficientKAN adopted the same approach as Liu et al. [1] but reworked the computation using B-spline basis functions followed by linear combination, reducing memory cost and simplifying computation [^4] https://github.com/Blealtan/efficient-kan. The authors replaced the incompatible L1 regularization on input samples with L1 regularization on weights. They also added learnable scales for activation functions and switched the base weight and spline scaler matrices to Kaiming uniform initialization, significantly improving performance on MNIST.
FastKAN can speed up training over EfficientKAN by using GRBFs to approximate the 3-order B-spline and employing layer normalization to keep inputs within the RBFs’ domain [5]. These modifications simplify the implementation without sacrificing accuracy. The GRBF has the formula:
$\phi(r)=e^{-\epsilon r^{2}}$ (7)
where $r=\|\mathbf{x}-\mathbf{c}\|$ is the Euclidean distance between the input vector $\mathbf{x}$ and the center vector $\mathbf{c}$, and $\epsilon$ ($epsilon>0$) is a sharp parameter that controls the width of the Gaussian function. FastKAN uses a special form of RBFs where $r=\frac{\mathbf{x}-\mathbf{c}}{h}$ and $\epsilon=\frac{1}{2}$, we can rewrite Equation 7 as [5]:
$\phi(r)=\exp\left(-\frac{1}{2}\left(\frac{\mathbf{x}-\mathbf{c}}{h}\right)^{2}\right)$ (8)
which $h$ controls the width of the Gaussian function. Finally, the RBF network with $N$ centers can be shown as [5]:
$RBF(\mathbf{x})=\sum_{i=1}^{N}w_{i}\phi(r_{i})=\sum_{i=1}^{N}w_{i}\exp\left(-\frac{1}{2}\left(\frac{\mathbf{x}-\mathbf{c}_{i}}{h}\right)^{2}\right)$ (9)
where $w_{i}$ represents adjustable weights or coefficients, and $\phi$ denotes the radial basis function as in Equation 7.
FasterKAN outperforms FastKAN in both forward and backward processing speeds [19]. It uses Reflectional Switch Activation Functions (RSWAFs), which are variants of RBFs and straightforward to compute due to their uniform grid structure. The RSWAF function $\phi(r)$ is shown as:
$\phi(r)=1-\left(\tanh\left(\frac{\mathbf{x}-\mathbf{c}}{h}\right)\right)^{2}$ (10)
Then, the RSWAF network with $N$ centers will be:
[…]
GottliebKAN, which is based on Gottlieb polynomials, showed the best result compared to the other 17 polynomial basic functions for setting up KANs on MNIST [23]. Gottlieb polynomials are a family of polynomials that arise in the study of the Bernoulli numbers and have applications in combinatorics and number theory [26]. Teymoor Seydi [23] denoted Gottlieb polynomials as:
$G_{n}(\mathbf{x})=\sum_{k=0}^{n}\binom{n}{k}B_{k}(\mathbf{x})\mathbf{x}^{n-k}$ (12)
where $G_{n}(\mathbf{x})$ represents the $n$-th Gottlieb polynomial, $(n\;k)$ denotes the binomial coefficient "n choose k", $B_{k}(\mathbf{x})$ are the Bernstein basis polynomials. Finally, $\mathbf{x}^{n-k}$ is the power term that adjusts the degree of $B_{k}(\mathbf{x})$ within the polynomial $G_{n}(\mathbf{x})$.
3.4 BSRBF-KAN
Inspired by successful implementations of previous KAN models, we developed BSRBF-KAN by combining B-splines and RBFs, which are widely recognized in computational mathematics for interpolation and approximation tasks. The RBFs we chose here are Gaussian ones because they are popular. Both B-splines and RBFs are smooth functions, ensuring continuity and derivatives up to a specified order, making them suitable for choosing proper $\Phi_{q}$ and $\phi_{q,p}$ in KANs.
From a variable $x$ in the vector input $\mathbf{x}$, the BSRBF function $\phi(x)$ is represented as:
$\phi(x)=w_{b}b(x)+w_{s}(\phi_{BS}(x)+\phi_{RBF}(x))$ (13)
where $b(x)$ and $w_{b}$ are the base output and its base matrix. $\phi_{BS}$ and $\phi_{RBF}$ are B-Spline and RBF functions, and $w_{s}$ is the spline matrix. Given the goal of designing a combined KAN, our intuition understands that the forward and backward speeds will not surpass those of individual KANs.
4 Experiments
Figure 1: The logarithm values of training losses (left) and validation losses (right) in 15 epochs in a training run.
We conducted 5 training runs for each model on MNIST [6] and Fashion-MNIST [7] to assess their overall performance more reliably. By taking the best metric value from all runs, we aimed to mitigate the effects of variability in training and ensure that the evaluation reflects the models’ peak capabilities. With the intention of designing simple networks, we only use activation functions, linear transformation, and layer normalization.
All models have a network structure of (784, 64, 10), comprising 784 input neurons, 64 hidden neurons, and 10 output neurons corresponding to the 10 output classes (0-9). For GottliebKAN, we used a structure (784, 64, 64, 10) to make it the same as the original design [^5] https://github.com/seydi1370/Basis_Functions. We trained the models with 15 epochs on MNIST and 25 epochs on Fashion-MNIST. Other hyperparameters are the same in all models, including batch_size=64, learning_rate=1e-3, weight_decay=1e-4, gamma=0.8, optimize=AdamW, and loss=CrossEntropy.
Figure 1 demonstrates training losses of models represented by logarithm values in a certain training run. On both MNIST and Fashion-MNIST, the training loss values of BSRBF-KAN drop sharply continuously, which makes it the model with the smallest loss after the training process. While this accelerates the training, there is also a concern that the model might be overfitting. We examined both training and validation accuracies across datasets and found no anomalies. In contrast, GottliebKAN lost the least loss, which affected its performance, as shown in the next experiments.
Table 1: The best metric values in 5 training runs on MNIST and Fashion-MNIST.
Dataset Model Train. Acc. Val. Acc. F1 Time (seconds) #Params
MNIST BSRBF-KAN 100.0 97.63 97.6 222 459040
FastKAN 99.94 97.38 97.34 102 459114
FasterKAN 98.52 97.38 97.36 93 408224
EfficientKAN 99.34 97.54 97.5 122 508160
GottliebKAN 99.66 97.78 97.74 269 219927
MLP 99.42 97.69 97.66 273 52512
Fashion-MNIST BSRBF-KAN 99.3 89.59 89.54 219 459040
FastKAN 98.27 89.62 89.6 160 459114
FasterKAN 94.4 89.39 89.3 157 408224
EfficientKAN 94.83 89.11 89.04 182 508160
GottliebKAN 93.79 87.69 87.61 241 219927
MLP 93.58 88.51 88.44 147 52512
Average of MNIST + Fashion-MNIST BSRBF-KAN 99.65 93.61 93.57 220.5 459040
FastKAN 99.11 93.50 93.47 131 459114
FasterKAN 96.46 93.39 93.33 125 408224
EfficientKAN 97.09 93.33 93.27 152 508160
GottliebKAN 96.73 92.74 92.68 255 219927
MLP 96.50 93.10 93.05 210 52512
Train. Acc = Training Accuracy, Val. Acc. = Validation Accuracy, #Params = Parameters
Table 2: The average metric values in 5 training runs on MNIST and Fashion-MNIST.
Dataset Model Train. Acc. Val. Acc. F1 Time (seconds)
MNIST BSRBF-KAN 100.00 ± 0.00 97.55 ± 0.03 97.51 ± 0.03 231
FastKAN 99.94 ± 0.01 97.25 ± 0.03 97.21 ± 0.03 101
FasterKAN 98.48 ± 0.01 97.28 ± 0.06 97.25 ± 0.06 93
EfficientKAN 99.37 ± 0.04 97.37 ± 0.07 97.33 ± 0.07 120
GottliebKAN 98.44 ± 0.61 97.19 ± 0.22 97.14 ± 0.23 221
MLP 99.44 ± 0.01 97.62 ± 0.03 97.59 ± 0.03 181
Fashion-MNIST BSRBF-KAN 99.19 ± 0.03 89.33 ± 0.07 89.29 ± 0.07 211
FastKAN 98.19 ± 0.04 89.42 ± 0.07 89.38 ± 0.07 162
FasterKAN 94.40 ± 0.01 89.26 ± 0.06 89.17 ± 0.07 154
EfficientKAN 94.76 ± 0.06 88.92 ± 0.08 88.85 ± 0.09 183
GottliebKAN 90.66 ± 1.08 87.16 ± 0.24 87.07 ± 0.25 238
MLP 93.56 ± 0.05 88.39 ± 0.06 88.36 ± 0.05 148
Average of MNIST + Fashion-MNIST BSRBF-KAN 99.60 93.44 93.40 221
FastKAN 99.07 93.34 93.30 131.5
FasterKAN 96.44 93.27 93.21 123.5
EfficientKAN 97.07 93.15 93.09 151.5
GottliebKAN 94.55 92.18 92.11 229.5
MLP 96.50 93.01 92.98 164.5
Train. Acc = Training Accuracy, Val. Acc. = Validation Accuracy
4.1 MNIST results analysis
As shown in Table 1, GottliebKAN is better than the other models, with a validation accuracy of 97.78% and an F1 score of 97.74%. We assume this because of its deeper network structure. MLP and BSRBF-KAN achieved second-best and third-best performances but took a long time to train. Conversely, FasterKAN had the shortest training time, but its performance slightly surpassed FastKAN. BSRBF-KAN is the only model with a training accuracy of 100% after 15 epochs. If counting the results based on average values, MLP is the best model, as displayed in Table 2, followed competitively by BSRBF-KAN. Despite achieving the highest accuracy values, GottliebKAN is unstable.
4.2 Fashion-MNIST results analysis
The ease of achieving high accuracy of models trained on MNIST may not accurately reflect their actual performance. Thus, we experimented with the model training on Fashion-MNIST. As shown in Table 1 and Table 2, FastKAN obtained the highest validation accuracy. At the same time, BSRBF-KAN followed competitively with and still kept the best convergence ability. MLP is the second-worst model, only performing better than GottliebKAN, but it has the fastest training time and the fewest parameters. However, when considering training time, we see that FasterKAN’s training time is only slightly longer than MLP’s, while achieving nearly 0.9% higher validation accuracy, according to the results in Table 1.
4.3 MNIST + Fashion-MNIST results analysis
When calculating the average values on MNIST and Fashion-MNIST, we clearly see that BSRBF-KAN achieved the best performance and also the best convergence, according to Table 1 and Table 2. Furthermore, all KANs except for GottliebKAN surpassed MLP. This indicates the positive signal of using KANs to obtain better accuracy values in model training, at least in these experiments.
4.4 Ablation Study
Several ablation experiments were performed to help us understand the impact of various components within BSRBF-KAN on overall model performance. Table 3 presents the metric values for different combinations of BSRBF-KAN components. Initially, we trained the model with all components included, followed by training variants with some components removed. The results indicate that the base output and layer normalization are key for the model’s performance, as their both absence leads to the worst results. The base output
While B-splines and RBFs can enhance performance, removing RBFs have less impact on the model than B-splines. Interestingly, when both are absent, the model becomes an MLP, which exhibits slightly better validation accuracy and F1 scores than the complete BSRBF-KAN model on MNIST, but on Fashion-MNIST, this is pretty bad. Thus, we believe that retaining B-splines and RBFs remains beneficial for the model’s performance.
Table 3: The performance of BSRBF-KAN by different components on MNIST and Fashion-MNIST in a training run.
Dataset Components Train. Acc. Val. Acc. F1
MNIST Full 100.0 97.53 97.49
No BS - 0.01 - 0.21 - 0.21
No RBF - 0.003 - 0.14 - 0.14
No BS + No RBF = MLP - 0.55 + 0.05 + 0.06
No BO - 0.01 - 0.41 - 0.41
No LN - 1.7 - 1.5 - 1.52
No BO + No LN - 5.16 - 5.97 - 6.08
Fashion-MNIST Full 99.39 89.43 89.4
No BS - 0.16 - 0.39 - 0.37
No RBF - 0.34 - 0.14 - 0.15
No BS + No RBF = MLP -5.86 -0.81 -0.78
No BO - 0.04 - 0.26 - 0.29
No LN - 6.12 - 1.1 - 1.13
No BO + No LN -6.23 -3.29 -3.38
Train. Acc = Training Accuracy, Val. Acc. = Validation Accuracy
No BS = No B-spline, No RBF = No Radial Basis Function
No BO = No Base Output, No LN = No Layer Normalization
5 Conclusion
We introduced BSRBF-KAN, a novel KAN that integrates B-splines and RBFs to fit input data in the training process. We did not intend to make BSRBF-KAN better than MLP and other KANs by optimizing hyperparameters; instead, we built it by default to understand aspects of KANs combination in training models. In the experiments, we trained BSRBF-KAN, MLP, and other KANs on the MNIST and Fashion-MNIST datasets. BSRBF-KAN demonstrated competitive performance and rapid convergence compared to other models. Furthermore, it is the best model if counting on the average values of both datasets. However, its high convergence rate can lead to overfitting, which can be mitigated by using dropout and other techniques.
All models were generally stable, except for GottliebKAN. Although GottliebKAN achieved the highest validation accuracy on MNIST, it performed the worst on Fashion-MNIST. Similarly, while MLP achieved the highest average validation accuracy on MNIST, it underperformed compared to other KANs on Fashion-MNIST. Therefore, it is crucial to experiment with our approach on other datasets in the future to gain a more comprehensive understanding. We will also continue to explore the combination of KANs and their components when designing new KAN architectures to enhance model performance across a wider range of problems.